The RMC Corporation blends three raw materials to produce two products: a fuel additive and a solvent base. Each ton of fuelMin 1 di s.t. 15 Material 1 0 x1 3 Material 2 8 3 3 30 Material 3 4 00 + 32 Goal 1 O X2 1 x2 -1 dit -1 dat X1, X2, dit, diidaC. Choose the correct graph that solves the model formulated in part b. (i) 90 x2 90 4x2 Material 3 80 Material 3 80 70+ 70 6(iii) (iv) 90 x2 90 4x2 Material 3 80 Material 3 80 70- 70+ 60- 60- Tons of Solvent Base 50- Tons of Solvent Base 50 40+ 40+(i) (ii) 90 4x2 90 4x2 Material 3 Material 3 80 80 70- 70 60+ 60+ Tons of Solvent Base 50- Tons of Solvent Base 50 40 40+ Mat(iii) (iv) 90 #x2 Material 3 80 90 x2 Material 3 80 70+ 70+ 60 60+ Tons of Solvent Base 50 - Tons of Solvent Base 50+ (12, 24everything is correct I just need the answer to part D on which graph is right.
Answer #1

Sol:

The Given data is tabulated as follows:

Resources

Requirement of each resource per ton

Available

Fuel Additive

Solvent Base

RM 1 (in ton)

2/5

1/2

20

RM 2 (in ton)

0

1/5

5

RM 3 (in ton)

3/5

3/10

21

a.

Amount needed to produce x1 = 30 and x2 = 15

Available amount

Comment

Raw Material

1

(2/5)(30) + (1/2)(15)

= 12 + 7.5= 19.5

20

Requirement <Available

2

(0)(30) + (1/5)(15) = 3

5

Requirement <Available

3

(3/5)(30) + (3/10)(15) = 18 +4.5 = 22.5

21

Requirement >Available

To satisfy both the goals, raw material 3 required is 22.5 units, but available is only 21 units, thus it is not possible to achieve both goals. Any of the goal will be underachieved.

b.

Min

__1__d1-

+

__1__d2-

s.t.

__2/5__x1

+

_1/2__x2

20

_____Material 1

_0__x1

+

_1/5___x2

5

______Material 2

__3/5__x1

+

__3/10__x2

21

_____Material 3

__1__x1

+

__0__x2

+

__1__d1+

+

__1__d1-

=

30

______Goal 1

__0__x1

+

__1___x2

+

___−1__d2+

+

__1__d2-

=

15

_______Goal 2

c.

Graph

Extreme point

d1-

d2-

Objective function Value

Z = d1- + d2-

i

(18.75, 25)

30 – 17.75 = 11.25

0

11.25

ii

(30, 10)

30-30 = 0

15 – 10 = 5

5

iii

(27.5, 15)

30-27.5 = 2.5

15-15 = 0

2.5

iv

(25, 20)

5

0

5

At the point (27.5, 15) the objective function value is minimum.

The objective function is minimum at point (27.5, 15), graph (iii) is correct graph, optimal product mix is:

fuel additive = 27.5 tons

solvent base = 15 tons

Correct Option: Graph (iii)

d.

If the goal 1 is twice important a goal 2, new objective function is Min: P1d1- + P2d2-

P1 = first priority to achieve at least 30 units of x1

P2 = second priority to achieve at least 15 units of x2

First identify point at which d1-is minimum, from four points the d1-is minimum at (30, 10).

Graph

Extreme point

d1-

d2-

Objective function Value

Z = d1- + d2-

i

(18.75, 25)

30 – 17.75 = 11.25

0

11.25

ii

(30, 10)

30-30 = 0

15 – 10 = 5

5

iii

(27.5, 15)

30-27.5 = 2.5

15-15 = 0

2.5

iv

(25, 20)

5

0

5

Thus, optimal product mix is:

fuel additive = 30 tons

solvent base = 10 tons

Correct Option: Graph (ii)

If you Satisfy with Answer, Please give me "Thumb Up". It is very useful for me.
Thank you for your support.